The sum of the area of two squares is 89 square centimeters. The length of a side of the larger square is 3 cm more than the length of a side of the smaller square. find the dimensions of each square.

Answer by sachi(548)

The length of a side of the larger square is 3 cm more than the length of a side of the smaller square.
let the side of the smaller square is a cm then larger square has sides a+3 cm
The sum of the area of two squares is 89 square centimeters.
so a^2+(a+3)^2=89
or a^2+a^2+9+6a=89
or 2a^2+6a-80=0
or a^2+3a-40=0
or a^2+8a-5a-40=0
or a(a+8)-5(a+8)=0
or (a+8)(a-5)=0
or a+8=0=> a=-8 rejected for -ve value
then (a-5=0 =>a=5 the side of the smaller square
& the sides of the larger square is 3+5=8 cms
ans